OpenStudy (anonymous):

A manufacturer determines that the number of drills it can sell is given by the formula D=-3p^2+180p-280, where p is the price of the drills in dollars.

OpenStudy (anonymous):

a. At what price will the manufacturer sell the maximum number of drills? b. What is the maximum number of drills that can be sold?

OpenStudy (anonymous):

Max point happens at x = -b/2a

OpenStudy (anonymous):

Here x is p!

OpenStudy (anonymous):

Understood but i am still confused....

OpenStudy (anonymous):

on this question.... i keep trying to get a. but it always turns out to be 84. here are the answer choices $30; 2,420 drills $90; 8,380 drills $60; 280 drills $32; 2,427 drills

OpenStudy (anonymous):

What do you get with p = 180/ 6 = ...?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

u can use differentiation to obtain the results on differentiating u have -6p+180=0 or p=30 at this value of p the above expression may be maximum or minimum to check u differentiate (the result once more) now it stands -6 thus at p=30 the second ordr differentiation is neagtive and hence at p=30, D is amximum in order to obtain the max value of D use p=30 in D hope rest u get it...

OpenStudy (anonymous):

Okay i got $30; 2,420 drills Is this correct for both questions?

OpenStudy (anonymous):

Does it take you 16hrs to plug 30 into the equation ?

OpenStudy (anonymous):

no, im sorry i fell asleep.